if it rains, an umbrella salesman can earn 20$ per day .if it's fair day he loses 5$ per day .What is his expectations if the probability of rain is 0.3?
Answer by ikleyn(52803)

.

His expectation is  20*0.3 + (-5)*(1-0.3) = 20*0.3 - 5*0.7 = 2.5 dollars.